# N. Deng & Y. Wang & W. Cao

### v1

@@ Conclusion
In this work, we extend the classical ergodicity theory of Da Prato & Debussche (1996) for the stochastic Cahn–Hilliard equation (SCHE) from the negative Sobolev space $\dot{H}^{-1}_{\alpha}$ to the physically more relevant Hilbert space $H_{\alpha}$, and we show that uniqueness of invariant measures fails on $H=L^{2}(\mathcal{O})$. We further introduce a fully discrete scheme—finite differences in space combined with a tamed exponential Euler method in time—for which uniform-in-time moment bounds are established. This enables us to prove uniform strong convergence with rates $1-\varepsilon$ in space and $\frac{3}{8}-\frac{\varepsilon}{2}$ in time.
@@
Author Contributions
Nan Deng: Formal alanalysis, Software, Visualization, Writing – Original Draft. Yibo Wang: Conceptualization, Methodology, Writing – Review &\& Editing. Wanrong Cao: Conceptualization, Methodology, Formal analysis, Writing – Review &\& Editing, Supervision, Funding acquisition.
Data Availability
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Statements and Declarations
This work was supported by the National Natural Science Foundation of China (No. 12571420), and the funding from the Jiangsu Provincial Scientific Research Center of Applied Mathematics (No. BK20233002). The authors have no relevant financial or non-financial interests to disclose.
@@
Proof 6.2.
Observe $\lambda_{0}=0$ and $\lambda_{j}\geq 1$ for $j\geq 1$. It follows from expansion in terms of the eigenbasis of $A$, Parseval’s identity and (3) that
@@
which yields (4). Similarly, we can obtain
@@
which together with (3) yields
@@
It follows from the orthonormality of $\left\{\phi_{j}\right\}_{j\geq 0}$ and (3) with $\alpha=\frac{2\alpha_{1}+5}{4}$ that
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\mathcal{L}_{1}^{N}(t,x,y)\right|^{2}\mathrm{d}y\leq\sup_{x\in\mathcal{O}}\sum_{j=N}^{\infty}\lambda_{j}^{2}e^{-2\lambda_{j}^{2}t}|\phi_{j}(x)|^{2}\leq Ce^{-\lambda_{1}^{2}t}t^{\frac{2\alpha_{1}+5}{4}}\sum_{j=N}^{\infty}\lambda_{j}^{-\frac{2\alpha_{1}+1}{2}}\leq Ce^{-\lambda_{1}^{2}t}t^{\frac{2\alpha_{1}+5}{4}}N^{-2\alpha_{1}}.$ (90)
We recall Eq. (3.10) from Lemma 3.2 in Cardon-Weber (2000), which states
@@
This identity, together with (3) with $\alpha=\frac{2\alpha_{1}+5}{4}$, implies
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\mathcal{L}_{2}^{N}(t,x,y)\right|^{2}\mathrm{d}y=$ $\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\sum_{j=1}^{N-1}\lambda_{N,j}^{2}e^{-2\lambda_{N,j}^{2}t}\left|\bar{\phi}_{N,j}(x)\right|^{2}\left|\left(\phi_{j}\left(\kappa_{N}(y)\right)-\phi_{j}(y)\right)\right|^{2}\mathrm{d}y$ (91)
@@
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\mathcal{L}_{3}^{N}(t,x,y)\right|^{2}\mathrm{d}y=$ $\sum_{j=1}^{N-1}\lambda_{N,j}^{2}e^{-2\lambda_{N,j}^{2}t}\sup_{x\in\mathcal{O}}\left|\phi_{j}(x)-\bar{\phi}_{N,j}(x)\right|^{2}$ (92)
@@
and applying (3) with $\alpha=\frac{6\alpha_{1}+5}{4}$ and (3) with $\hat{\alpha}=\frac{\alpha_{1}}{2}$, we obtain
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\mathcal{L}_{4}^{N}(t,x,y)\right|^{2}\mathrm{d}y=$ $\sum_{j=1}^{N-1}\lambda_{N,j}^{2}e^{-2\lambda_{N,j}^{2}t}\left(1-e^{-2\left(\lambda_{j}^{2}-\lambda_{N,j}^{2}\right)t}\right)^{2}\leq Ce^{-\lambda_{N,1}^{2}t}t^{-\frac{2\alpha_{1}+5}{4}}N^{-2\alpha_{1}}.$ (93)
@@
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\mathcal{L}_{5}^{N}(t,x,y)\right|^{2}\mathrm{d}y=$ $\sum_{j=1}^{N-1}\left(\lambda_{j}-\lambda_{N,j}\right)^{2}e^{-2\lambda_{j}^{2}t}\leq Ce^{-\lambda_{N,1}^{2}t}t^{-\frac{2\alpha_{1}+5}{4}}N^{-2\alpha_{1}}.$ (94)
By Hölder’s inequality and (90)–(94) one can get
@@
This completes the proof.
@@
$\tilde{u}^{N}(t,x)=$ $\int_{\mathcal{O}}G_{t}^{N}(x,y)u_{0}\left(\kappa_{N}(y)\right)\mathrm{~d}y+\int_{0}^{t}\int_{\mathcal{O}}\Delta_{N}G_{t-s}^{N}(x,y)f\left(u\left(s,\kappa_{N}(y)\right)\right)\mathrm{~d}y\mathrm{d}s$ (95)
@@
$\left\{\begin{aligned} &\mathrm{d}\tilde{U}^{N}_{t}+A_{N}^{2}\tilde{U}^{N}_{t}\mathrm{d}t=A_{N}F_{N}(\mathbb{U}_{t})\mathrm{d}t+\sigma\sum_{j=1}^{N-1}\phi_{N,j}\mathrm{d}\beta_{j}(t),\quad t>0,\\ &\tilde{U}_{0}=U_{0}.\end{aligned}\right$. (96)
@@
$\sup_{t\geq 0}\left\|u(t,\cdot)-\tilde{u}^{N}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq Ch^{1-\epsilon}.$ (97)
Following the proof of Deng& Cao (2025)[Proposition 3.3], we can directly verify that
@@
For any $x\in\mathcal{O}$, according to the definitions of $\tilde{u}^{N}_{0}$ and $\Delta_{N}$ and Assumption 2, we have
$|\tilde{u}^{N}_{0}(x)-u_{0}(x)|\leq Ch^{2}\ \text{and}\ |\Delta_{N}u_{0}(\kappa_{N}(x))-u^{\prime\prime}_{0}(x)|\leq Ch^{2}.$ (98)
It follows from (98), (44) and (57) with $\alpha_{1}=1-\epsilon$ that
$\left\|\mathcal{I}_{1}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq\sup_{x\in\mathcal{O}}\big|\tilde{u}^{N}_{0}(x)-u_{0}(x)\big|\leq Ch^{2},$ (99)
@@
$\left\|\mathcal{I}_{2}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq$ $\int_{0}^{t}\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\Delta_{N}G_{t-s}^{N}(x,y)-\Delta G_{t-s}(x,y)\right|\left(\|f\left(u\left(s,\kappa_{N}(y)\right)\right)\|_{L^{p}(\Omega;\mathbb{R})}+\Delta_{N}u_{0}(y)\right)\mathrm{d}y\mathrm{d}s$ (100)
@@
To estimate of $\mathcal{I}_{3}(t,x)$, we need to investigate $u\left(t,\kappa_{N}(x)\right)-u(t,x)$. Taking an arbitrarily small positive number $\epsilon$, we can choose a sufficiently large number $q$ such that $q>\max\left\{p+2,3/\epsilon\right\}$. For any $t>0$ and $x\in\mathcal{O}$, utilizing Assumption 2, (61), Hölder’s inequality, (57) and (3) with $\alpha=\frac{15}{8}$, we obtain
@@
$\left\|\mathcal{I}_{3}(t,x)\right\|_{L^{q}(\Omega;\mathbb{R})}$ (101)
@@
Moreover, for any $x,\ z\in\mathcal{O}$, by (12) with $\alpha=\frac{3}{4}$ we obtain
$\left\|\mathcal{I}_{3}(t,x)-\mathcal{I}_{3}(t,z)\right\|_{L^{q}(\Omega;\mathbb{R})}\leq Ch\int_{0}^{t}\int_{\mathcal{O}}\left|\Delta G_{s}(x,y)-\Delta G_{s}(z,y)\right|\mathrm{d}y\mathrm{d}s\leq Ch\int_{0}^{t}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left|x-z\right|^{\frac{3}{4}}\mathrm{d}s\leq Ch|x-z|^{\frac{3}{4}}.$
Thus we employ Revuz & Yor (1999, Theorem 2.1) to obtain
@@
which combined with (101) yields
$\left\|\mathcal{I}_{3}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq\left\|\mathcal{I}_{3}(t,\cdot)\right\|_{L^{q}(\Omega;L^{\infty})}\leq\left\|\mathcal{I}_{3}(t,\cdot)-\mathcal{I}_{3}(t,1)\right\|_{L^{q}(\Omega;L^{\infty})}+\left\|\mathcal{I}_{3}(t,1)\right\|_{L^{q}(\Omega;\mathbb{R})}\leq Ch.$ (102)
@@
$\left\|\mathcal{I}_{4}(t,x)\right\|_{L^{q}(\Omega;\mathbb{R})}^{2}\leq C\sum_{k=1}^{3}\mathcal{I}_{4,k}(t,x),$ (103)
@@
A proof similar to the one used in (90), (92), and (93) shows that
$\left\|\mathcal{I}_{4}(t,x)\right\|_{L^{q}(\Omega;\mathbb{R})}\leq C\sqrt{\sum_{k=1}^{3}\mathcal{I}_{4,k}(t,x)}\leq Ch^{1-\frac{\epsilon}{2}}.$ (104)
According to (11) and (61), one has
@@
Combining this with Revuz & Yor (1999, Theorem 2.1) leads to
$\mathbb{E}\left[\sup_{x\neq z}\left(\frac{|\mathcal{I}_{4}(t,x)-\mathcal{I}_{4}(t,z)|}{|x-z|^{1-\frac{2}{q}}}\right)^{q}\right]\leq C.$ (105)
Then it can follow from (104) and (105) that
$\left\|\mathcal{I}_{4}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq$ $\left\|\mathcal{I}_{4}(t,\cdot)\right\|_{L^{q}(\Omega;L^{\infty})}$ (106)
@@
$\sup_{t\geq 0}\left\|\tilde{u}^{N}(t,\cdot)-u^{N}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq Ch^{1-\epsilon}.$ (107)
@@
$\left\{\begin{aligned} &\mathrm{d}E_{t}^{N}+A_{N}^{2}E_{t}^{N}\mathrm{d}t=A_{N}\left[F_{N}(U_{t}^{N}+E_{t}^{N}+(\mathbb{U}_{t}-\tilde{U}^{N}_{t}))-F_{N}(U_{t}^{N})\right]\mathrm{d}t,\ t\in(0,\infty),\\ &E_{0}^{N}=0.\end{aligned}\right$. (108)
@@
$\|E_{t}^{N}\|_{L^{p}(\Omega;l^{2}_{N})}\leq$ $C\sqrt{K_{5}(t)}\left(\mathbb{I}_{1}\big(t,\frac{1}{2},\frac{\lambda_{N,1}^{2}}{2},\big\|\mathbb{U}-\tilde{U}^{N}\big\|_{L^{2p}(\Omega;l^{2}_{N}))}^{2}\big)\right)^{\frac{1}{2}}$ (109)
@@
By utilizing (52) and Lemma 4.4 one can get that
$\left\|\tilde{U}_{t}^{N}\right\|_{w^{1,2}_{N}}$ $\leq\left\|e^{-A_{N}^{2}t}u^{N}_{0}\right\|_{w^{1,2}_{N}}+\int_{0}^{t}\left\|\left(-A_{N}\right)^{\frac{3}{2}}e^{-A_{N}^{2}(t-s)}F_{N}\left(\mathbb{U}_{s}\right)\right\|_{l_{N}^{2}}\mathrm{d}s+\left\|\mathscr{O}_{t}^{N}\right\|_{w^{1,2}_{N}}$ (110)
@@
This proof is completed.
@@
For the base case $m=1$, it follows from (55) and the orthonormality of $\left\{\phi_{N,j}\right\}_{j=0}^{N-1}$ that
@@
which, combined with the fact that $\{\beta_{0,j}\}_{j=1}^{N-1}$ are Gaussian random variables with mean zero, confirms that (111) holds.
Assume now that (111) holds for $m=k$. For $m=k+1$, it again follows from (55) and orthonormality that
$\langle U^{\tau,N}_{t_{k+1}},\phi_{N,j}\rangle_{l^{2}_{N}}=$ $e^{-\lambda_{N,j}^{2}\tau}\langle U^{\tau,N}_{t_{k}},\phi_{N,j}\rangle_{l^{2}_{N}}+\tau\lambda_{N,j}e^{-\lambda_{N,j}^{2}\tau}\langle\tilde{F}(U^{\tau,N}_{t_{k}}),\phi_{N,j}\rangle_{l^{2}_{N}}+\sigma e^{-\lambda_{N,j}^{2}\tau}\langle\Delta\beta_{k},\phi_{N,j}\rangle_{l^{2}_{N}}$ (112)
@@
$\langle\tilde{F}(U^{\tau,N}_{t_{k}}),\phi_{N,j}\rangle_{l^{2}_{N}}=$ $\frac{1}{1+\tau\big\|U^{\tau,N}_{t_{k}}\big\|_{w^{1,2}_{N}}^{12}}\Big\langle a_{0}\Big(U^{\tau,N}_{t_{k}}\Big)^{3}+a_{1}\Big(U^{\tau,N}_{t_{k}}\Big)^{2}+a_{2}U^{\tau,N}_{t_{k}}+a_{3},\phi_{N,j}\Big\rangle_{l^{2}_{N}}$ (113)
@@
By substituting (113) into (112) and using the symmetry of $\{\tilde{X}_{k,j}\}_{j=1}^{N-1}$ and $\{\beta_{k,j}\}_{j=1}^{N-1}$, we deduce that (111) also holds for $m=k+1$. Thus, by the principle of mathematical induction, the claim holds for all $m\in\mathbb{N}$.
It then follows that $\mathbb{E}\left[\phi_{v,1,\alpha_{2}}(u^{\tau,N}(t_{m},\cdot))\right]=0$ for all $m\in\mathbb{N}$. From (83), we conclude that
@@
which completes the proof.

### v3

@@ Conclusion
In this work, we extend the classical ergodicity theory of [15] for the stochastic Cahn–Hilliard equation (SCHE) from the negative Sobolev space $\dot{H}^{-1}_{\alpha}$ to the physically more relevant Hilbert space $H_{\alpha}$, and we show that uniqueness of invariant measures fails on $H=L^{2}(\mathcal{O})$. We further introduce a fully discrete scheme—finite differences in space combined with a tamed exponential Euler method in time—for which uniform-in-time moment bounds are established. This enables us to prove uniform strong convergence with rates $1-\varepsilon$ in space and $\frac{3}{8}-\frac{\varepsilon}{2}$ in time.
@@
Proof.
Observe $\lambda_{0}=0$ and $\lambda_{j}\geq 1$ for $j\geq 1$. It follows from expansion in terms of the eigenbasis of $A$, Parseval’s identity and (2.2) that
@@
which yields (2.3). Similarly, we can obtain
@@
which together with (2.2) yields
@@
It follows from the orthonormality of $\left\{\phi_{j}\right\}_{j\geq 0}$ and (2.2) with $\alpha=\frac{2\alpha_{1}+5}{4}$ that
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\mathcal{L}_{1}^{N}(t,x,y)\right|^{2}\mathrm{d}y\leq\sup_{x\in\mathcal{O}}\sum_{j=N}^{\infty}\lambda_{j}^{2}e^{-2\lambda_{j}^{2}t}|\phi_{j}(x)|^{2}\leq Ce^{-\lambda_{1}^{2}t}t^{\frac{2\alpha_{1}+5}{4}}\sum_{j=N}^{\infty}\lambda_{j}^{-\frac{2\alpha_{1}+1}{2}}\leq Ce^{-\lambda_{1}^{2}t}t^{\frac{2\alpha_{1}+5}{4}}N^{-2\alpha_{1}}.$ (6.2)
We recall Eq. (3.10) from Lemma 3.2 in [7], which states
@@
This identity, together with (2.2) with $\alpha=\frac{2\alpha_{1}+5}{4}$, implies
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\mathcal{L}_{2}^{N}(t,x,y)\right|^{2}\mathrm{d}y=$ $\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\sum_{j=1}^{N-1}\lambda_{N,j}^{2}e^{-2\lambda_{N,j}^{2}t}\left|\bar{\phi}_{N,j}(x)\right|^{2}\left|\left(\phi_{j}\left(\kappa_{N}(y)\right)-\phi_{j}(y)\right)\right|^{2}\mathrm{d}y$ (6.3)
@@
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\mathcal{L}_{3}^{N}(t,x,y)\right|^{2}\mathrm{d}y=$ $\sum_{j=1}^{N-1}\lambda_{N,j}^{2}e^{-2\lambda_{N,j}^{2}t}\sup_{x\in\mathcal{O}}\left|\phi_{j}(x)-\bar{\phi}_{N,j}(x)\right|^{2}$ (6.4)
@@
and applying (2.2) with $\alpha=\frac{6\alpha_{1}+5}{4}$ and (2.2) with $\hat{\alpha}=\frac{\alpha_{1}}{2}$, we obtain
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\mathcal{L}_{4}^{N}(t,x,y)\right|^{2}\mathrm{d}y=$ $\sum_{j=1}^{N-1}\lambda_{N,j}^{2}e^{-2\lambda_{N,j}^{2}t}\left(1-e^{-2\left(\lambda_{j}^{2}-\lambda_{N,j}^{2}\right)t}\right)^{2}\leq Ce^{-\lambda_{N,1}^{2}t}t^{-\frac{2\alpha_{1}+5}{4}}N^{-2\alpha_{1}}.$ (6.5)
@@
$\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\mathcal{L}_{5}^{N}(t,x,y)\right|^{2}\mathrm{d}y=$ $\sum_{j=1}^{N-1}\left(\lambda_{j}-\lambda_{N,j}\right)^{2}e^{-2\lambda_{j}^{2}t}\leq Ce^{-\lambda_{N,1}^{2}t}t^{-\frac{2\alpha_{1}+5}{4}}N^{-2\alpha_{1}}.$ (6.6)
By Hölder’s inequality and (6.2)–(6.6) one can get
@@
This completes the proof. ∎
@@
$\tilde{u}^{N}(t,x)=$ $\int_{\mathcal{O}}G_{t}^{N}(x,y)u_{0}\left(\kappa_{N}(y)\right)\mathrm{~d}y+\int_{0}^{t}\int_{\mathcal{O}}\Delta_{N}G_{t-s}^{N}(x,y)f\left(u\left(s,\kappa_{N}(y)\right)\right)\mathrm{~d}y\mathrm{d}s$ (6.7)
@@
$\left\{\begin{aligned} &\mathrm{d}\tilde{U}^{N}_{t}+A_{N}^{2}\tilde{U}^{N}_{t}\mathrm{d}t=A_{N}F_{N}(\mathbb{U}_{t})\mathrm{d}t+\sigma\sum_{j=1}^{N-1}\phi_{N,j}\mathrm{d}\beta_{j}(t),\quad t>0,\\ &\tilde{U}_{0}=U_{0}.\end{aligned}\right$. (6.8)
@@
$\sup_{t\geq 0}\left\|u(t,\cdot)-\tilde{u}^{N}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq Ch^{1-\epsilon}.$ (6.9)
Following the proof of [19][Proposition 3.3], we can directly verify that
@@
For any $x\in\mathcal{O}$, according to the definitions of $\tilde{u}^{N}_{0}$ and $\Delta_{N}$ and Assumption 2.3, we have
$|\tilde{u}^{N}_{0}(x)-u_{0}(x)|\leq Ch^{2}\ \text{and}\ |\Delta_{N}u_{0}(\kappa_{N}(x))-u^{\prime\prime}_{0}(x)|\leq Ch^{2}.$ (6.10)
It follows from (6.10), (3.36) and (4.11) with $\alpha_{1}=1-\epsilon$ that
$\left\|\mathcal{I}_{1}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq\sup_{x\in\mathcal{O}}\big|\tilde{u}^{N}_{0}(x)-u_{0}(x)\big|\leq Ch^{2},$ (6.11)
@@
$\left\|\mathcal{I}_{2}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq$ $\int_{0}^{t}\sup_{x\in\mathcal{O}}\int_{\mathcal{O}}\left|\Delta_{N}G_{t-s}^{N}(x,y)-\Delta G_{t-s}(x,y)\right|\left(\|f\left(u\left(s,\kappa_{N}(y)\right)\right)\|_{L^{p}(\Omega;\mathbb{R})}+\Delta_{N}u_{0}(y)\right)\mathrm{d}y\mathrm{d}s$ (6.12)
@@
To estimate of $\mathcal{I}_{3}(t,x)$, we need to investigate $u\left(t,\kappa_{N}(x)\right)-u(t,x)$. Taking an arbitrarily small positive number $\epsilon$, we can choose a sufficiently large number $q$ such that $q>\max\left\{p+2,3/\epsilon\right\}$. For any $t>0$ and $x\in\mathcal{O}$, utilizing Assumption 2.3, (4.15), Hölder’s inequality, (4.11) and (2.2) with $\alpha=\frac{15}{8}$, we obtain
@@
$\left\|\mathcal{I}_{3}(t,x)\right\|_{L^{q}(\Omega;\mathbb{R})}$ (6.13)
@@
Moreover, for any $x,\ z\in\mathcal{O}$, by (3.4) with $\alpha=\frac{3}{4}$ we obtain
$\left\|\mathcal{I}_{3}(t,x)-\mathcal{I}_{3}(t,z)\right\|_{L^{q}(\Omega;\mathbb{R})}$ $\leq Ch\int_{0}^{t}\int_{\mathcal{O}}\left|\Delta G_{s}(x,y)-\Delta G_{s}(z,y)\right|\mathrm{d}y\mathrm{d}s$
$\leq Ch\int_{0}^{t}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left|x-z\right|^{\frac{3}{4}}\mathrm{d}s\leq Ch|x-z|^{\frac{3}{4}}.$
Thus we employ [34, Theorem 2.1] to obtain
@@
which combined with (6.13) yields
$\left\|\mathcal{I}_{3}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq\left\|\mathcal{I}_{3}(t,\cdot)\right\|_{L^{q}(\Omega;L^{\infty})}\leq\left\|\mathcal{I}_{3}(t,\cdot)-\mathcal{I}_{3}(t,1)\right\|_{L^{q}(\Omega;L^{\infty})}+\left\|\mathcal{I}_{3}(t,1)\right\|_{L^{q}(\Omega;\mathbb{R})}\leq Ch.$ (6.14)
@@
$\left\|\mathcal{I}_{4}(t,x)\right\|_{L^{q}(\Omega;\mathbb{R})}^{2}\leq C\sum_{k=1}^{3}\mathcal{I}_{4,k}(t,x),$ (6.15)
@@
A proof similar to the one used in (6.2), (6.4), and (6.5) shows that
$\left\|\mathcal{I}_{4}(t,x)\right\|_{L^{q}(\Omega;\mathbb{R})}\leq C\sqrt{\sum_{k=1}^{3}\mathcal{I}_{4,k}(t,x)}\leq Ch^{1-\frac{\epsilon}{2}}.$ (6.16)
According to (3.3) and (4.15), one has
@@
Combining this with [34, Theorem 2.1] leads to
$\mathbb{E}\left[\sup_{x\neq z}\left(\frac{|\mathcal{I}_{4}(t,x)-\mathcal{I}_{4}(t,z)|}{|x-z|^{1-\frac{2}{q}}}\right)^{q}\right]\leq C.$ (6.17)
Then it can follow from (6.16) and (6.17) that
$\left\|\mathcal{I}_{4}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq$ $\left\|\mathcal{I}_{4}(t,\cdot)\right\|_{L^{q}(\Omega;L^{\infty})}$ (6.18)
@@
$\sup_{t\geq 0}\left\|\tilde{u}^{N}(t,\cdot)-u^{N}(t,\cdot)\right\|_{L^{p}(\Omega;L^{\infty})}\leq Ch^{1-\epsilon}.$ (6.19)
@@
$\left\{\begin{aligned} &\mathrm{d}E_{t}^{N}+A_{N}^{2}E_{t}^{N}\mathrm{d}t=A_{N}\left[F_{N}(U_{t}^{N}+E_{t}^{N}+(\mathbb{U}_{t}-\tilde{U}^{N}_{t}))-F_{N}(U_{t}^{N})\right]\mathrm{d}t,\ t\in(0,\infty),\\ &E_{0}^{N}=0.\end{aligned}\right$. (6.20)
@@
$\|E_{t}^{N}\|_{L^{p}(\Omega;l^{2}_{N})}\leq$ $C\sqrt{K_{5}(t)}\left(\mathbb{I}_{1}\big(t,\frac{1}{2},\frac{\lambda_{N,1}^{2}}{2},\big\|\mathbb{U}-\tilde{U}^{N}\big\|_{L^{2p}(\Omega;l^{2}_{N}))}^{2}\big)\right)^{\frac{1}{2}}$ (6.21)
@@
By utilizing (4.6) and Lemma 4.4 one can get that
$\left\|\tilde{U}_{t}^{N}\right\|_{w^{1,2}_{N}}$ $\leq\left\|e^{-A_{N}^{2}t}u^{N}_{0}\right\|_{w^{1,2}_{N}}+\int_{0}^{t}\left\|\left(-A_{N}\right)^{\frac{3}{2}}e^{-A_{N}^{2}(t-s)}F_{N}\left(\mathbb{U}_{s}\right)\right\|_{l_{N}^{2}}\mathrm{d}s+\left\|\mathscr{O}_{t}^{N}\right\|_{w^{1,2}_{N}}$ (6.22)
@@
This proof is completed. ∎
@@
For the base case $m=1$, it follows from (4.9) and the orthonormality of $\left\{\phi_{N,j}\right\}_{j=0}^{N-1}$ that
@@
which, combined with the fact that $\{\beta_{0,j}\}_{j=1}^{N-1}$ are Gaussian random variables with mean zero, confirms that (6.23) holds.
Assume now that (6.23) holds for $m=k$. For $m=k+1$, it again follows from (4.9) and orthonormality that
$\langle U^{\tau,N}_{t_{k+1}},\phi_{N,j}\rangle_{l^{2}_{N}}=$ $e^{-\lambda_{N,j}^{2}\tau}\langle U^{\tau,N}_{t_{k}},\phi_{N,j}\rangle_{l^{2}_{N}}+\tau\lambda_{N,j}e^{-\lambda_{N,j}^{2}\tau}\langle\tilde{F}(U^{\tau,N}_{t_{k}}),\phi_{N,j}\rangle_{l^{2}_{N}}+\sigma e^{-\lambda_{N,j}^{2}\tau}\langle\Delta\beta_{k},\phi_{N,j}\rangle_{l^{2}_{N}}$ (6.24)
@@
$\langle\tilde{F}(U^{\tau,N}_{t_{k}}),\phi_{N,j}\rangle_{l^{2}_{N}}=$ $\frac{1}{1+\tau\big\|U^{\tau,N}_{t_{k}}\big\|_{w^{1,2}_{N}}^{12}}\Big\langle a_{0}\Big(U^{\tau,N}_{t_{k}}\Big)^{3}+a_{1}\Big(U^{\tau,N}_{t_{k}}\Big)^{2}+a_{2}U^{\tau,N}_{t_{k}}+a_{3},\phi_{N,j}\Big\rangle_{l^{2}_{N}}$ (6.25)
@@
By substituting (6.25) into (6.24) and using the symmetry of $\{\tilde{X}_{k,j}\}_{j=1}^{N-1}$ and $\{\beta_{k,j}\}_{j=1}^{N-1}$, we deduce that (6.23) also holds for $m=k+1$. Thus, by the principle of mathematical induction, the claim holds for all $m\in\mathbb{N}$.
It then follows that $\mathbb{E}\left[\phi_{v,1,\alpha_{2}}(u^{\tau,N}(t_{m},\cdot))\right]=0$ for all $m\in\mathbb{N}$. From (5.4), we conclude that
@@
which completes the proof. ∎